# Probing Quantum Spin Systems with Kolmogorov-Arnold Neural Network Quantum States

### v1

@@ Abstract
Neural Quantum States (NQS) are a class of variational wave functions parametrized by neural networks (NNs) to study quantum many-body systems. In this work, we propose SineKAN, the NQS ansatz based on Kolmogorov-Arnold Networks (KANs), to represent quantum mechanical wave functions as nested univariate functions. We show that SineKAN wavefunction with learnable sinusoidal activation functions can capture the ground state energies, fidelities and various correlation functions of the 1D Transverse-Field Ising model, Anisotropic Heisenberg model, and Antiferromagnetic $J_{1}-J_{2}$ model with different chain lengths. In our study of the $J_{1}-J_{2}$ model with $L=100$ sites, we find that the SineKAN model outperforms several previously explored neural quantum state ansätze, including Restricted Boltzmann Machines (RBMs), Long Short-Term Memory models (LSTMs), and Feed-Forward Neural Networks (FFCN), when compared to the results obtained from the Density Matrix Renormalization Group (DMRG) algorithm.
@@ II Kolmogorov-Arnold Variational Wavefunction
III Benchmark Neural Networks
In section VI.1, we benchmark the SineKAN ansätze against three neural networks: MLP, RNN-LSTM, and RBM.
Previous work has been done on the use of RNNs to find GSs and their associated wave functions [6, 7]. The motivation behind the choice of RNN architectures is that they generally encode some information about long-term and short-term “memory” across the length of inputs in a sequence. The idea in this case is that the memory of the models could capture short- and/or long-range dependencies across the spin lattice. Here we use a long-short-term memory (RNN-LSTM) as compared to alternative GRU or simple RNN used in [6, 7] due to previous literature findings that RNN-LSTM generally outperform the other RNN models on large datasets and sequences where very-long-term memory is not necessary [39, 40].
Another well studied ansatz for NQS is the RBM [3, 41]. The strongest motivation for the RBM model is high computational efficiency for smaller input sizes allowing scaling to much larger sizes compared to RNN-LSTM or SineKAN at comparable computation times for lattices on the order of $L=100$. RBM models are effectively classical Ising models that have a connected graph equivalent to the graph of an input lattice structure. Due to this, they are expected to perform well in NQS particularly for minimally frustrated systems.
@@ III Benchmark Neural Networks
RBM ([41, 43]) are also known as the restricted stochastic Ising-Lenz-Little model. They are constructed from neurons arranged in a bipartite graph so that all layers are generally separable. There is a strong physics motivation to compare these because they are directly analogous to the transverse field Ising model. The computational complexity of RBMS is more comparable to $\mathcal{O}(e^{N})$ so they are highly efficient when lattices sizes remain relatively small.
@@ VII.3 Benchmark of $J_{1}-J_{2}$ Model Against ED
For $J_{2}>J_{2}^{c}$, the system becomes gapped, with a dimerized two-fold degenerate GS that breaks translational invariance [55]. The emergence of this dimer order can be probed using the connected dimer–dimer correlation function: [16, 56]
@@ Acknowledgments
Appendix A Appendix: Variational Monte Carlo

### v2

@@ Abstract
Neural Quantum States (NQS) are a class of variational wave functions parametrized by neural networks (NNs) to study quantum many-body systems. In this work, we propose SineKAN, a NQS ansatz based on Kolmogorov-Arnold Networks (KANs), to represent quantum mechanical wave functions as nested univariate functions. We show that SineKAN wavefunction with learnable sinusoidal activation functions can capture the ground state energies, fidelities and various correlation functions of the one dimensional Transverse-Field Ising model, Anisotropic Heisenberg model, and Antiferromagnetic $J_{1}-J_{2}$ model with different chain lengths. In our study of the $J_{1}-J_{2}$ model with $L=100$ sites, we find that the SineKAN model outperforms several previously explored neural quantum state ansätze, including Restricted Boltzmann Machines (RBMs), Long Short-Term Memory models (LSTMs), and Feed-Forward Neural Networks (FFNN), when compared to the results obtained from the Density Matrix Renormalization Group (DMRG) algorithm. We find that SineKAN models can be trained to high precisions and accuracies with minimal computational costs.
@@ II Kolmogorov-Arnold Variational Wavefunction
III Benchmark NQS Models
We evaluate the effectiveness of the SineKAN ansätze by benchmarking against three widely used NQS baselines: the RBM, RNN-LSTM, and MLP, as described in section VI.1.
As the original NQS proposed by Carleo and Troyer, RBM has served as a foundational benchmark in variational studies of quantum many-body systems [3, 39]. Its main advantage is that it’s high computational efficiency for small to medium-sized systems, which allows it to scale to larger systems—up to around 100 spins—with reasonable training time, especially compared to more complex architectures like RNN-LSTM. The computational complexity of RBM is more comparable to $\mathcal{O}(e^{N})$, so they remain highly efficient as long as the lattice size is relatively small. Structurally, the RBM consists of a visible layer representing spin configurations and a hidden layer of binary units, connected through trainable weights without intra-layer couplings ([39, 40]). Due to this bipartite structure, RBM is particularly effective in NQS applications for unfrustrated or weakly frustrated systems.
The application of RNNs to represent GS wavefunctions in quantum spin systems is well established, with several studies demonstrating their effectiveness [6, 7]. The motivation behind the choice of RNN architectures is that they generally encode some information about long-term and short-term “memory” across the length of inputs in a sequence. The idea in this case is that the memory of the models could capture short- and/or long-range dependencies across the spin lattice. Here we use a long-short-term memory (RNN-LSTM) as compared to alternative GRU or simple RNN used in [6, 7] due to previous literature findings that RNN-LSTM generally outperform the other RNN models on large datasets and sequences where very-long-term memory is not necessary [41, 42].
@@ VII.3 Benchmark of $J_{1}-J_{2}$ Model Against ED
For $J_{2}>J_{2}^{c}$, the system becomes gapped, with a dimerized two-fold degenerate GS that breaks translational invariance [55]. The emergence of this dimer order can be probed using the connected dimer–dimer correlation function: [16, Becca3]
@@ Acknowledgments
Appendix A Variational Monte Carlo
@@
In practice, the sampling of configurations $\sigma$ is typically performed using the Metropolis–Hastings algorithm, which generates a Markov chain of spin configurations distributed according to the probability $p_{\theta}(\sigma)$. Once a sufficiently large and equilibrated set of samples is collected, the gradient estimates can be computed. For the optimization step, simple stochastic gradient descent may be used, but more sophisticated techniques such as the stochastic reconfiguration (SR) [2, 64] method—also known as natural gradient descent—are often preferred. These take into account the underlying geometry of the variational manifold, leading to improved stability and convergence properties in high-dimensional parameter spaces.